OpenStudy (iwanttomakearequest):

Transform the following polar equation into an equation in rectangular coordinates: r=-4 sin (theta)

OpenStudy (iwanttomakearequest):

options are: A) y= -4 B) x + y= -4 C) x^2 + (y+2)^2 = 4 D) x = -4y

OpenStudy (.whitedragon.):

sorry wrong one

OpenStudy (iwanttomakearequest):

no worries :)

OpenStudy (.whitedragon.):

you can still use this We multiply both sides by R R = 4 sin t R 2 = 4 R sin t We now use the relationship between polar and rectangular coordinates: R 2 = x 2 + y 2 and y = R sin t

OpenStudy (iwanttomakearequest):

Is that R^2 = x^2 + y^2 or just R 2 = x 2 + y 2?

OpenStudy (.whitedragon.):

yes ^ its R^2 x^2 and Y^2

OpenStudy (iwanttomakearequest):

okay, so what's x and what's y??

OpenStudy (.whitedragon.):

i dont know i looked up the answer heres where i got it http://www.analyzemath.com/polarcoordinates/polar_to_rectangular_eq.html

OpenStudy (iwanttomakearequest):

oh okay thanks, I'll take a look at that. But it looks like it's solving for 4sin(theta) not -4sin(theta)

OpenStudy (.whitedragon.):

yeah i noticed that after i posted it sorry thats the best I can do

OpenStudy (iwanttomakearequest):

All right, thanks anyway :)

OpenStudy (.whitedragon.):

yup